In a certain dart game, a player can score either 1 for landing in the outer ring, 3 for landing in the middle ring, and 5 for landing in the inner ring. Assume a player makes 60 throws, and assume they have an equal chance of hitting each ring. What is the chance the player scores at least 200 points?
Answer #1
x f(x) xP(x) x2P(x)
1 1/3 0.333 0.3333
3 1/3 1.000 3.0000
5 1/3 1.667 8.3333
total 3.0000 11.6667
E(x) =μ= ΣxP(x) = 3.0000
E(x2) = Σx2P(x) = 11.6667
Var(x)=σ2 = E(x2)-(E(x))2= 2.6667
std deviation=         σ= √σ2 = 1.6330

expected score for 60 throws =60*3 =180

and standard deviation =1.633*sqrt(60)= 12.649

therefore from Normal approximation:

chance the player scores at least 200 points:

P(X>200)=P(Z>(200-180)/12.649)=P(Z>1.58)=0.0569